# RIZE: Regularized Imitation Learning via Distributional Reinforcement Learning

### v1

@@ 6   Conclusion
We present a novel IRL framework that overcomes the rigidity of fixed reward mechanisms through dynamic reward adaptation via context-sensitive regularization and probabilistic return estimation with distributional RL. By introducing adaptive target rewards that evolve during training—replacing static assignments—our method enables nuanced alignment of expert and agent trajectories, while value distribution integration captures richer return dynamics without sacrificing theoretical consistency. Empirical evaluations on MuJoCo benchmarks demonstrate state-of-the-art performance, achieving expert-level proficiency on the Humanoid task with only three demonstrations, alongside ablation studies confirming the necessity of applying our regularization mechanism. This work advances imitation learning by unifying flexible reward learning with probabilistic return representations, offering a scalable, sample-efficient paradigm for complex decision making. Future directions include extending these principles to offline IL and risk-sensitive robotics, where adaptive rewards and uncertainty-aware distributions could further enhance robustness and generalization.
Appendix A Proofs
@@
Consider an MDP with discount factor $\gamma\in[0,1)$ and rewards bounded by some constant $C>0$ such that $|R(s,a)|\leq C$. Define the return distribution as
$Z(s,a)=\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t}),$
where the trajectory $\{(s_{t},a_{t})\}$ is generated by following policy $\pi$ under dynamics $\mathcal{P}$. Then the classical action-value function is given by
$Q^{\pi}(s,a)=\mathbb{E}_{\pi,\mathcal{P}}[Z(s,a)].$
Moreover, if we approximate $Z(s,a)$ by a learned distribution $Z_{\theta}(s,a)$ with expectation $\tilde{Q}(s,a)=\mathbb{E}[Z_{\theta}(s,a)]$, then under the conditions of the Dominated Convergence Theorem (DCT), $\tilde{Q}(s,a)$ converges to $Q^{\pi}(s,a)$.
Proof of Lemma A.1.
Existence of the Expectation: Since the rewards are bounded, we have $|R(s_{t},a_{t})|\leq C$, and the discounted series satisfies
$\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}|R(s_{t},a_{t})|\right]\leq\sum_{t=0}^{\infty}\gamma^{t}C=\frac{C}{1-\gamma}<\infty.$
This absolute convergence allows us to apply DCT:
$\mathbb{E}[Z(s,a)]=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}[R(s_{t},a_{t})].$
Equivalence to Q-value: By definition,
$Q^{\pi}(s,a)=\mathbb{E}_{\pi,\mathcal{P}}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]=E[Z(s,a)].$
Convergence of the Learned Distribution: In distributional reinforcement learning, if we represent the return distribution using $Z_{\theta}(s,a)$, then its mean is given by
$\tilde{Q}(s,a)=E[Z_{\theta}(s,a)].$
Under DCT conditions, we have
$\tilde{Q}(s,a)\to Q^{\pi}(s,a).$
Proof of Lemma 4.1.
1. Bounded Rewards
The expert and policy reward targets $\lambda^{\pi_{E}},\lambda^{\pi}$ are learned via:
$\lambda^{\pi_{E}}\leftarrow\mathbb{E}_{\rho_{E}}[R_{Q}],\quad\lambda^{\pi}\leftarrow\mathbb{E}_{\rho_{\pi}}[R_{Q}].$
This ensures $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$ track the empirical mean of $R_{Q}$.
Fixing $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$, solving $\nabla_{R_{Q}}\Gamma=0$ yields:
$R_{Q}=\frac{\rho_{E}\lambda^{\pi_{E}}+\rho_{\pi}\lambda^{\pi}}{\rho_{E}+\rho_{\pi}},$
This is a convex combination of $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$. Since $\lambda^{\pi_{E}},\lambda^{\pi}$ are empirical means, $|R_{Q}|\leq\max(|\lambda^{\pi_{E}}|,|\lambda^{\pi}|)$, ensuring boundedness.
2. Temporal Consistency
The reward $R_{Q}$ (for either expert or policy samples) is defined via the soft Q-function:
$R_{Q}=Q(s,a)-\gamma\mathbb{E}_{s^{\prime}}\left[Q(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})\right].$
Rearranging gives:
$Q(s,a)=R_{Q}+\gamma\mathbb{E}_{s^{\prime}}\left[Q(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})\right]+\epsilon,$
where $\epsilon$ is the Bellman optimization error (due to approximation or finite samples). Substituting $R_{Q}=\lambda$:
$Q(s,a)-\gamma Q(s^{\prime},a^{\prime})=\lambda+\epsilon-\gamma\alpha\log\pi(a^{\prime}|s^{\prime}).$
Taking absolute values:
$|Q(s,a)-\gamma Q(s^{\prime},a^{\prime})|\leq|\lambda|+|\epsilon|+\gamma\alpha|\log\pi(a^{\prime}|s^{\prime})|.$
This bounds successive Q-values, ensuring temporal consistency.
Conclusion.
The adaptive regularizer $\Gamma$ (Equation 5) enforces bounded rewards via convex combinations of empirical targets and ensures temporal consistency through entropy-regularized Q-value recursion. Both properties stabilize learning by preventing divergence. ∎
Appendix B Experiments
@@ B.1   MuJoCo Control Suite
We evaluate our imitation learning approach, RIZE, on four benchmark Gym [Brockman et al., 2016] MuJoCo [Todorov et al., 2012] environments: HalfCheetah-v2, Walker2d-v2, Ant-v2, and Humanoid-v2. Expert trajectories are generated using a pretrained Soft Actor-Critic (SAC) [Haarnoja et al., 2018] agent, with each trajectory consisting of 1,000 state-action transitions. To facilitate performance comparisons, episode returns are normalized relative to expert performance: HalfCheetah (5100), Walker2d (5200), Ant (4700), and Humanoid (5300).
@@ B.2   Implementation Details
Our architecture integrates components from Distributional SAC (DSAC) [^2] https://github.com/xtma/dsac [Ma et al., 2020] and IQ-Learn [^3] https://github.com/Div99/IQ-Learn [Garg et al., 2021], with hyperparameters tuned through systematic search and ablation studies. Key configurations for experiments involving three demonstrations are summarized in Table 1, while Table 2 provides settings for scenarios with ten demonstrations. Our implementation can be found at https://github.com/adibka/RIZE.
Distributional SAC Components: The critic network is implemented as a three-layer multilayer perceptron (MLP) with 256 units per layer, trained using a learning rate of $3\times 10^{-4}$. The policy network is a four-layer MLP, also with 256 units per layer. Environment-specific learning rates are applied: $1\times 10^{-5}$ for the Humanoid environment and $5\times 10^{-5}$ for all others. To enhance training stability, we employ a target policy—a delayed version of the online policy—and sample next-state actions from this module. For value distribution training $Z^{\pi}_{\phi,\tau}$, we adopt the Implicit Quantile Networks (IQN) [Dabney et al., 2018a] approach by sampling quantile fractions $\tau$ uniformly from $\mathcal{U}(0,1)$. Additionally, dual critic networks with delayed updates are used, which empirically improve training stability.
@@
    $\mathcal{L}(\pi,Q)=\mathbb{E}_{\rho_{E}}[Q^{\pi}(s,a)-\gamma V^{\pi}(s^{\prime})]-\mathbb{E}_{\rho}[V^{\pi}(s)-\gamma V^{\pi}(s^{\prime})]-c\Gamma(R_{Q},\lambda)$ (9)
  2. 2.
@@
    $\mathcal{L}(\pi,Q)=\mathbb{E}_{\rho_{E}}[Q^{\pi}(s,a)-\gamma V^{\pi}(s^{\prime})]-(1-\gamma)\mathbb{E}_{p_{0}}[V^{\pi}(s_{0})]-c\Gamma(R_{Q},\lambda)$ (10)
Here, $\rho$ is a mixture distribution, $p_{0}$ denotes the initial distribution, $R_{Q}(s,a)$ is the implicit reward defined as $R_{Q}(s,a)=Q^{\pi}(s,a)-\gamma V^{\pi}(s^{\prime})$, the state-value function is given by $V^{\pi}(s^{\prime})=Q^{\pi}(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})$, and lastly, our convex regularizer is expressed as $\Gamma(R_{Q},\lambda)=\mathbb{E}_{\rho_{E}}[(R_{Q}-\lambda^{\pi_{E}})^{2}]+\mathbb{E}_{\rho_{\pi}}[(R_{Q}-\lambda^{\pi})^{2}]$.
The choice between v0 or value loss variants depends on environment complexity: we find that for a complex task like Humanoid-v2, the v0 variant demonstrates greater robustness. And, for simpler tasks such as HalfCheetah-v2, Walker2d-v2, and Ant-v2, the value variant performs better.
Table 1: Three Demonstrations Hyperparameters
Environment $\alpha$ $c$ lr $\pi$ $\lambda^{\pi_{E}}$ lr $\lambda^{\pi_{E}}$ $\lambda^{\pi}$ lr $\lambda^{\pi_{E}}$
Ant-v2 0.05 0.1 $5\times 10^{-5}$ 10 $1\times 10^{-4}$ 5 $1\times 10^{-5}$
HalfCheetah-v2 0.05 0.1 $5\times 10^{-5}$ 10 $1\times 10^{-4}$ 5 $1\times 10^{-5}$
Walker2d-v2 0.05 0.1 $5\times 10^{-5}$ 10 $1\times 10^{-4}$ 5 $1\times 10^{-5}$
Humanoid-v2 0.05 0.5 $1\times 10^{-5}$ 10 $1\times 10^{-4}$ 5 $5\times 10^{-5}$
Table 2: 10 Demonstrations Hyperparameters
Environment $\alpha$ $c$ lr $\pi$ $\lambda^{\pi_{E}}$ lr $\lambda^{\pi_{E}}$ $\lambda^{\pi}$ lr $\lambda^{\pi_{E}}$
Ant-v2 0.1 0.1 $5\times 10^{-5}$ 10 $1\times 10^{-4}$ 5 $1\times 10^{-4}$
HalfCheetah-v2 0.1 0.1 $5\times 10^{-5}$ 10 $1\times 10^{-4}$ 5 $1\times 10^{-4}$
Walker2d-v2 0.1 0.1 $5\times 10^{-5}$ 10 $1\times 10^{-4}$ 5 $1\times 10^{-4}$
Humanoid-v2 0.1 0.5 $1\times 10^{-5}$ 10 $1\times 10^{-4}$ 5 $1\times 10^{-5}$
B.3   Extended Main Results
In this section, we showcase the performance of our method, RIZE, using both ten and three expert demonstrations. RIZE outshines LSIQ and SQIL in various tasks, with IQ-Learn as its only competitor. Remarkably, it reaches expert-level performance in the challenging Humanoid-v2 environment with just three trajectories, proving its scalability. Plus, it requires fewer training steps than other algorithms, emphasizing its efficiency and robustness in complex situations. Our findings demonstrate that using fewer expert samples does not negatively impact our imitation learning approach (see Figure 5).
B.4   Extended Recovered Rewards
Here, we investigate the implicit reward dynamics of our method using both three and ten expert demonstrations. As illustrated in Figure 6, with ample expert trajectories (n=10), the rewards for both the expert and the policy stabilize to consistent values across most environments. Notably, even with a three demonstrations, which presents a significant learning challenge, rewards still converge to similar values in tasks such as HalfCheetah and Humanoid.
Our implicit reward regularization term, $\Gamma(R_{Q},\lambda)$, ensures the learned rewards stay close to the target rewards. Looking at Figure X, you can see that optimizing $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$ using Equations X effectively bounds both the implicit and target rewards, boosting overall stability. Interestingly, for certain tasks like HalfCheetah-v2, Walker2d-v2, and Ant-v2, $\lambda^{\pi}$ doesn’t change much. This is because we deliberately used smaller learning rates, which proved crucial for maintaining training stability.
Figure 6: Implicit reward curves for expert and policy samples on Gym MuJoCo tasks. n represents the number of expert demonstrations.
Figure 7: Target reward values for expert ($\lambda^{\pi_{E}}$) and policy ($\lambda^{\pi}$) on Gym MuJoCo tasks. n represents the number of expert demonstrations.
B.5   Hyperparameter Fine-Tuning
@@
Selecting appropriate initial values and learning rates for the automatic fine-tuning of $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$ is critical in our approach. First, we observe that a suitable learning rate is essential for the stable training of our imitation learning agent, as illustrated in Figure 8(a). Our findings indicate that $\lambda^{\pi}$ must be optimized very slowly; using larger learning rates can destabilize training and hinder progress. In contrast, $\lambda^{\pi_{E}}$ demonstrates greater resilience when optimized with higher learning rates. Additionally, $\lambda^{\pi_{E}}$ remains robust even with varying initial values. However, as shown in Figure 8(b), failing to select an appropriate initial value for $\lambda^{\pi}$ can negatively impact learning. Overall, Figures 8(a) and 8(b) highlight the need for careful selection of both the learning rate and initial value when optimizing $\lambda^{\pi}$, while $\lambda^{\pi_{E}}$ exhibits considerable robustness in this regard.
Figure 8: Fine-tuning of target parameters. (a) Turquoise represents our method’s primary result with learning rates of $1e^{-4}$ for $\lambda^{\pi_{E}}$ and $1e^{-5}$ for $\lambda^{\pi}$. Orange and blue lines indicate higher learning rates (e.g., 0.001) for $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$, respectively. (b) Turquoise shows the main result with initial values of 10 for $\lambda^{\pi_{E}}$ and 5 for $\lambda^{\pi}$, while other lines explore different starting values. Only one parameter is varied at a time, and three trajectories are used throughout the analysis.
@@
Our experiments on the regularizer coefficient, $C$, reveal that smaller values of $C$ encourage expert-like performance, while larger values overly constrain rewards and targets, limiting learning. This finding highlights the critical role of selecting an appropriate $C$, as it directly impacts the balance between learning from expert data and regularization: higher values prioritize regularization at the cost of learning, whereas smaller values favor learning but reduce regularization (see Figure 9).
Figure 9: Fine-tuning of the regularizer coefficient ($C$). Turquoise shows our method’s primary result with $C=0.1$ for Ant-v2, while gray indicates a larger value ($C=0.6$). Only one parameter is varied at a time, and three trajectories are used throughout the analysis.
@@
We observe that the entropy coefficient is a crucial hyperparameter in inverse reinforcement learning (IRL) problems. As shown in Figure 10, IRL methods typically require small values for $\alpha$, a point previously noted by Reference X. With expert demonstrations available, an imitation learning (IL) policy does not need to explore for optimal actions, as these are provided by the demonstrations. Consequently, higher values of $\alpha$ can lead to training instability, ultimately resulting in policy collapse.
Figure 10: Fine-tuning of the temperature parameter ($\alpha$). Turquoise shows our method’s primary result with $\alpha=0.05$ for Ant-v2, while gray indicates a larger value ($\alpha=0.5$). Only one parameter is varied at a time, and three trajectories are used throughout the analysis.

### v2

@@ 6 Conclusion
We propose a novel IRL framework that overcomes the limitations of fixed reward mechanisms through dynamic reward adaptation and context-sensitive regularization. Our approach ensures bounded implicit rewards and stable value function updates, leading to robust policy optimization. By integrating distributional RL with implicit reward learning, we capture richer return dynamics while preserving theoretical guarantees. Empirical results on MuJoCo benchmarks show expert-like proficiency on the Humanoid task with only three demonstrations. Ablation studies confirm the critical role of our regularization mechanism. This work unifies implicit reward regularization with distributional return representations, offering a scalable and sample-efficient solution for complex decision-making. Future directions include extending this framework to offline IL and risk-sensitive robotics, where adaptive rewards and uncertainty-aware distributions can further improve robustness and generalization.
Appendix A Supporting Proofs
A.1 Expectation of the Value Distribution
@@
The expectation of the distributional return satisfies
$\mathbb{E}[Z(s,a)]=Q(s,a).$
Proof.
From the soft value distribution definition
$Z(s,a)=\sum_{t=0}^{\infty}\gamma^{t}\left[R(s_{t},a_{t})+\alpha\mathcal{H}(\pi(\cdot\mid s_{t}))\right]\,$, (14)
taking expectations yields:
$\mathbb{E}[Z(s,a)]$ $=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(R(s_{t},a_{t})+\alpha\mathcal{H}(\pi(\cdot\mid s_{t}))\right)\right]$
$=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}\left[R(s_{t},a_{t})+\alpha\mathcal{H}(\pi(\cdot\mid s_{t}))\right]$
$=Q(s,a)\quad\text{(by soft Q-value definition)}.$
A.2 Bounds on the Optimal Implicit Reward
Corollary A.2.
The optimal implicit reward satisfies:
$R_{Q}^{*}(s,a)\in\left[-\frac{1}{2c}+\lambda_{\min},\;\frac{1}{2c}+\lambda_{\max}\right]\,$, (15)
where $\lambda_{\min}:=\min\left\{\lambda^{\pi_{E}},\lambda^{\pi}\right\}$ and $\lambda_{\max}:=\max\left\{\lambda^{\pi_{E}},\lambda^{\pi}\right\}$.
Proof.
From Proposition 4.1:
$R_{Q}^{*}(s,a)=\underbrace{\frac{\rho_{E}-\rho_{\pi}}{2c(\rho_{E}+\rho_{\pi})}}_{\text{Term 1}}+\underbrace{\frac{\rho_{E}\lambda^{\pi_{E}}+\rho_{\pi}\lambda^{\pi}}{\rho_{E}+\rho_{\pi}}}_{\text{Term 2}}\,.$ (16)
Term 1 lies in $\left[-\frac{1}{2c},\;\frac{1}{2c}\right]$ since $\rho_{E},\rho_{\pi}\geq 0$ (achieved when either $\rho_{\pi}\to 0$ or $\rho_{E}\to 0$). Term 2 is a convex combination of $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$, thus in $\left[\lambda_{\min},\;\lambda_{\max}\right]$. Combining intervals yields the result. ∎
A.3 Convergence of the Optimal Reward
Corollary A.3.
For $\rho_{\pi}=\rho_{E}$, the optimal reward satisfies $R_{Q}^{*}(s,a)=\lambda^{\pi_{E}}=\lambda^{\pi}$.
Proof.
When $\rho_{\pi}=\rho_{E}$, Proposition 4.1 simplifies to
$R_{Q}^{*}(s,a)=\frac{\lambda^{\pi_{E}}+\lambda^{\pi}}{2}\,.$ (17)
Substituting this expression into the loss functions for $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$ in Equation (8) yields
$\underset{\lambda^{\pi_{E}}}{\min}\;\mathbb{E}_{\rho_{E}}\left[\frac{1}{4}\left(\lambda^{\pi}-\lambda^{\pi_{E}}\right)^{2}\right]\,,\qquad\underset{\lambda^{\pi}}{\min}\;\mathbb{E}_{\rho_{\pi}}\left[\frac{1}{4}\left(\lambda^{\pi_{E}}-\lambda^{\pi}\right)^{2}\right]\,.$ (18)
Differentiating each objective with respect to its corresponding target and setting the derivatives to zero gives $\lambda^{\pi}=\lambda^{\pi_{E}}$. This result implies that the optimal targets coincide at convergence. Therefore, the optimal reward reduces to
$R_{Q}^{*}(s,a)=\lambda^{\pi_{E}}=\lambda^{\pi}\,.$ (19)
Appendix B Additional Experiments
@@ B.1 MuJoCo Control Suite
We evaluate our imitation learning approach, RIZE, on four benchmark Gym (Brockman et al., 2016) MuJoCo (Todorov et al., 2012) environments: HalfCheetah-v2, Walker2d-v2, Ant-v2, and Humanoid-v2. Expert trajectories were obtained from IQ-Learn (Garg et al., 2021), generated by a Soft Actor-Critic agent (Haarnoja et al., 2018), with each trajectory comprising 1,000 state-action transitions. To facilitate performance comparisons, episode returns are normalized relative to expert performance: HalfCheetah (5100), Walker2d (5200), Ant (4700), and Humanoid (5300). All implementation details used in our experiments are publicly available at https://github.com/adibka/RIZE.
@@ B.2 Implementation Details
Our architecture integrates components from Distributional SAC (DSAC) [^2] https://github.com/xtma/dsac (Ma et al., 2020) and IQ-Learn [^3] https://github.com/Div99/IQ-Learn (Garg et al., 2021), with hyperparameters tuned through systematic search and ablation studies. Key configurations for experiments involving three and ten demonstrations are summarized in Table 1.
Distributional SAC Components: The critic network is implemented as a three-layer multilayer perceptron (MLP) with 256 units per layer, trained using a learning rate of $3\times 10^{-4}$. The policy network is a four-layer MLP, also with 256 units per layer. Environment-specific learning rates are applied: $1\times 10^{-5}$ for the Humanoid environment and $5\times 10^{-5}$ for all others. To enhance training stability, we employ a target policy—a delayed version of the online policy—and sample next-state actions from this module. For value distribution training $Z^{\pi}_{\phi,\tau}$, we adopt the Implicit Quantile Networks (IQN) (Dabney et al., 2018a) approach by sampling quantile fractions $\tau$ uniformly from $\mathcal{U}(0,1)$. Additionally, dual critic networks with delayed updates are used, which empirically improve training stability.
@@
    $\mathcal{L}(\pi,Q)=\mathbb{E}_{\rho_{E}}[Q(s,a)-\gamma V(s^{\prime})]-\mathbb{E}_{\rho}[V(s)-\gamma V(s^{\prime})]-c\,\Gamma(R_{Q},\lambda)$
@@
    $\mathcal{L}(\pi,Q)=\mathbb{E}_{\rho_{E}}[Q(s,a)-\gamma V(s^{\prime})]-(1-\gamma)\mathbb{E}_{p_{0}}[V(s_{0})]-c\,\Gamma(R_{Q},\lambda)$
Here, $\rho$ is a mixture distribution, $p_{0}$ denotes the initial distribution, $R_{Q}(s,a)$ is the implicit reward defined as $R_{Q}(s,a)=Q(s,a)-\gamma V(s^{\prime})$, the state-value function is given by $V(s^{\prime})=Q(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})$, and lastly, our convex regularizer is expressed as $\Gamma(R_{Q},\lambda)=\mathbb{E}_{\rho_{E}}[(R_{Q}-\lambda^{\pi_{E}})^{2}]+\mathbb{E}_{\rho_{\pi}}[(R_{Q}-\lambda^{\pi})^{2}]$.
The choice between v0 or value loss variants depends on environment complexity: we find that for a complex task like Humanoid-v2, the v0 variant demonstrates greater robustness. And, for HalfCheetah-v2, Walker2d-v2, and Ant-v2, the value variant performs better.
Table 1: Hyperparameters for 3 and 10 Demonstrations (merged where identical)
Environment $\alpha$ (3 / 10) $c$ lr $\pi$ lr $Z$ $\lambda^{\pi_{E}}$ lr $\lambda^{\pi_{E}}$ $\lambda^{\pi}$ lr $\lambda^{\pi}$ (3 / 10)
Ant-v2 0.05 / 0.10 0.1 $5\times 10^{-5}$ $3\times 10^{-4}$ 10 $1\times 10^{-4}$ 5 $\times 10^{-5} / 1×1$$\times 10^{-4}$
HalfCheetah-v2 0.05 / 0.10 0.1 $5\times 10^{-5}$ $3\times 10^{-4}$ 10 $1\times 10^{-4}$ 5 $\times 10^{-5} / 1×1$$\times 10^{-4}$
Walker2d-v2 0.05 / 0.10 0.1 $5\times 10^{-5}$ $3\times 10^{-4}$ 10 $1\times 10^{-4}$ 5 $\times 10^{-5} / 1×1$$\times 10^{-4}$
Humanoid-v2 0.05 / 0.10 0.5 $1\times 10^{-5}$ $3\times 10^{-4}$ 10 $1\times 10^{-4}$ 5 $5\times 10^{-5}$ / $1\times 10^{-5}$
B.3 Hyperparameter Tuning
@@
Selecting appropriate initial values and learning rates for the automatic fine-tuning of $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$ is critical in our approach. First, we observe that a suitable learning rate is essential for the stable training of our imitation learning agent, as illustrated in Figure 7(a). Our findings indicate that $\lambda^{\pi}$ must be optimized very slowly; using larger learning rates can destabilize training and hinder progress. In contrast, $\lambda^{\pi_{E}}$ demonstrates greater resilience when optimized with higher learning rates. Additionally, $\lambda^{\pi_{E}}$ remains robust even with varying initial values. However, as shown in Figure 7(b), failing to select an appropriate initial value for $\lambda^{\pi}$ can negatively impact learning. Overall, Figures 7(a) and 7(b) highlight the need for careful selection of both the learning rate and initial value when optimizing $\lambda^{\pi}$, while $\lambda^{\pi_{E}}$ exhibits considerable robustness in this regard.
Figure 7: Fine-tuning analysis of target parameters. (a) Learning rates: Turquoise represents our method’s primary result with learning rates of $1\mathrm{e}{-4}$ for $\lambda^{\pi_{E}}$ and $1\mathrm{e}{-5}$ for $\lambda^{\pi}$. Orange and blue lines indicate higher learning rates (e.g., $1\mathrm{e}{-3}$) for $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$, respectively. (b) Starting values: Turquoise shows the main result with initial values of 10 for $\lambda^{\pi_{E}}$ and 5 for $\lambda^{\pi}$, while other lines explore different starting values. Three trajectories are used throughout the analysis.
@@
Our experiments on the regularizer coefficient $c$ reveal that smaller values of $c$ encourage expert-like performance, while larger values overly constrain rewards and targets, limiting learning. This finding highlights the critical role of selecting an appropriate $c$, as it directly impacts the balance between learning from expert data and regularization: higher values prioritize regularization at the cost of learning, whereas smaller values favor learning but reduce regularization (see Figure 8(a)).
@@
We observe that the entropy coefficient is a crucial hyperparameter in inverse reinforcement learning (IRL) problems. As shown in Figure 8(b), IRL methods typically require small values for $\alpha$, a point previously noted (Garg et al., 2021). With expert demonstrations available, an imitation learning (IL) policy does not need to explore for optimal actions, as these are provided by the demonstrations. Consequently, higher values of $\alpha$ can lead to training instability, ultimately resulting in policy collapse.
Figure 8: (a) Effect of the regularizer coefficient $c$. Turquoise shows the primary result of our method with $c=0.1$, while gray represents a larger value ($c=0.6$). (b) Effect of the temperature parameter $\alpha$. Turquoise shows the result with $\alpha=0.05$, and gray corresponds to a larger value ($\alpha=0.5$). Three trajectories are used throughout the analysis.